# Contributions of $\rho(770,1450)\to\omega\pi$ for the Cabibbo-favored $D\to h\omega\pi$ decays

### v1

@@ Abstract
Recently, the BESIII Collaboration has observed three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\pi^{+}\omega$ and $D^{0}\to K^{-}\pi^{+}\omega$. In this work, we investigate the contributions of the subprocesses $\rho^{+}\to\omega\pi^{+}$ in these Cabibbo-favored decays $D\to h\omega\pi$, with $\rho^{+}=\{\rho(770)^{+},\rho(1450)^{+},\rho(770)^{+}\&\rho(1450)^{+}\}$ and $h=\{\eta,K^{0}_{S},K^{-}\}$, by introducing these subprocesses into the decay amplitudes of the relevant decay processes via the vector form factor $F_{\omega\pi}$ measured in the related $\tau$ and $e^{+}e^{-}$ processes; we provide the first theoretical predictions for the branching fractions of the quasi-two-body decays $D_{s}^{+}\to\eta[\rho^{+}\to]\omega\pi^{+}$, $D^{+}\to K^{0}_{S}[\rho^{+}\to]\omega\pi^{+}$ and $D^{0}\to K^{-}[\rho^{+}\to]\omega\pi^{+}$. Our findings reveal that the contributions from the subprocess $\rho(770)^{+}\to\omega\pi^{+}$ are significant in these observed three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\omega\pi^{+}$ and $D^{0}\to K^{-}\omega\pi^{+}$, notwithstanding the contributions originating from the Breit-Wigner tail effect of $\rho(770)^{+}$. The numerical results of this study suggest that these Cabibbo-favored three-body decays are dominated by the contributions that come from the $P$-wave intermediate states $\rho(770)^{+}$, $\rho(1450)^{+}$ and their interference effects.
@@ I Introduction
The three-body decay $D_{s}^{+}\to\eta\omega\pi^{+}$ has been observed for the first time by BESIII Collaboration recently, with its total branching fraction $\mathcal{B}=(0.54\pm 0.12_{\rm stat}\pm 0.04_{\rm syst})\%$ prd107.052010 . This result aligns an earlier measurement by the CLEO Collaboration, which reported a branching fraction of $(0.85\pm 0.54_{\rm stat}\pm 0.06_{\rm syst})\%$ in Ref. prd80.051102 for the same decay channel, but the BESIII measurement offers significantly improved precision. Given the quark structure of the initial and the final states in this $D_{s}$ decay process, the Cabibbo-favored transition $c\to s$ along with the $W^{+}\to u\bar{d}$ will be the dominated process at quark level. The prospective intermediate states for this process could be the resonances $a_{0}(980)^{+}$, $\rho(770)^{+}$, $b_{1}(1235)^{+}$, and $\omega(1420)$, etc., and their excited states which will decay into the $\pi^{+}\eta$, $\omega\pi^{+}$ and $\omega\eta$ pairs in the final state, respectively PDG-2024 . The combination of $\omega\eta$ and $\pi^{+}$ with the intermediate resonance $\omega(1420)$ means a pure annihilation process for this $D_{s}^{+}$ decay. The team of $\omega$ and $a_{0}(980)^{+}\to\pi^{+}\eta$ for this decay is very similar to the decay process $D_{s}^{+}\to\pi^{0}[a_{0}(980)^{+}\to]\pi^{+}\eta$, which has been measured by BESIII prl123.112001 and the branching fraction was found to be consistent with the triangle rescattering processes via the intermediate processes $D_{s}^{+}\to\eta^{(\prime)}\rho^{+}$, $D_{s}^{+}\to K^{*}\bar{K}$ and $D_{s}^{+}\to K\bar{K^{*}}$ epjc80-895 ; plb803-135279 ; epjc80-1041 ; prd103.116016 ; prd109.076027 ; epjc81-1093 . But the branching fraction for $D_{s}^{+}\to\omega a_{0}(980)^{+}$ was found to be less than $3.4\times 10^{-4}$ in Ref. epjc81-1093 , due to the cancellation of the rescattering effects and the suppressed short-distance $W$ annihilation contribution. Consequently, the dominated contributions for the decay $D_{s}^{+}\to\eta\omega\pi^{+}$ are expected to come from the combination of $\eta$ and the resonances which decay into $\omega\pi^{+}$ pair in the final state.
@@
The light meson pair of $\omega\pi$ in the three-body hadronic $D$ and $B$ meson decays, which originating from the weak current of the matrix element $\langle\omega\pi|V^{\mu}\!-\!A^{\mu}|0\rangle$, is related to the processes $\tau\to\omega\pi\nu_{\tau}$ and $e^{+}e^{-}\to\omega\pi$. The $G$-parity conservation requires that the $\omega\pi$ pair is determined mainly by the vector part $V^{\mu}$ of the weak current zpc57-495 , which is associated with the $\rho$ family resonances as the intermediate states via the $P$-wave transition amplitudes prd86.037302 . The axial-vector part $A^{\mu}$ of the weak current for $\omega\pi$ system goes to the resonance $b_{1}(1235)$ and its excited states. The state $b_{1}(1235)$, with the quantum numbers $J^{PG}=1^{++}$, is related to the second-class weak current classified in terms of parity and $G$ parity pr112.1375 . The second-class weak current has been searched in $\tau$ decays in recent years, but no evidence has been observed ar0807.4900 ; npb218-110 . The decay amplitude for the $b_{1}(1235)\to\omega\pi$ subprocess through $S$- and $D$-waves in relevant processes is proportional to the mass difference between $u$ and $d$ quarks PDG-2024 , which makes the contribution of $\omega\pi$ from the axial-vector resonance $b_{1}(1235)$ accompanied by the second-class current highly suppressed.
@@
The ordinary decay mode for $\rho(770)\to\omega\pi$ is not allowed because of the pole mass of $\rho(770)$, which is apparently below the threshold of the $\omega\pi$ pair PDG-2024 . But the Breit-Wigner (BW) BW-model tail effect of $\rho(770)$, also known as the virtual contribution plb25-294 ; Dalitz62 ; prd94.072001 ; plb791-342 , was found indispensable to the production of $\omega\pi$ pair in the processes of $\tau\to\omega\pi\nu_{\tau}$ plb185-223 ; prd61.072003 ; rmp78.1043 ; prl103.041802 and $e^{+}e^{-}\to\omega\pi^{0}$ plb174-453 ; plb486-29 ; plb562-173 ; plb669-223 ; JETPL94-734 ; prd88.054013 ; prd94.112001 ; prd96.092009 ; plb813-136059 ; 2309.00280 . The excited state $\rho(1450)$ has been observed to decay into $\omega\pi$ pair in $B$ meson decays by the CLEO, Babar and Belle collaborations prd64.092001 ; prd74.012001 ; prd92.012013 . This state has been suggested as a $2S$-hybrid mixture in Ref. prd60.114011 in view of its decay characters npb443-233 ; prd52.1706 ; prd56.1584 , but its mass is consistent with that for the $2S$ excitation of $\rho(770)$ prd55.4157 . The further investigation of the interference between the $\rho(1450)$ and its ground state would provide deeper insights into its nature.
@@
In addition to data for the $D_{s}^{+}\to\eta\omega\pi^{+}$ decay, the absolute branching fractions of the decays $D^{0}\to K^{-}\pi^{+}\omega$, $D^{0}\to K^{0}_{S}\pi^{0}\omega$, and $D^{+}\to K^{0}_{S}\pi^{+}\omega$ have been determined by BESIII Collaboration recently in Ref. prd105.032009 , with the results to be $(3.392\pm 0.044_{\rm stat}\pm 0.085_{\rm syst})\%$, $(0.848\pm 0.046_{\rm stat}\pm 0.031_{\rm syst})\%$, and $(0.707\pm 0.041_{\rm stat}\pm 0.029_{\rm syst})\%$, respectively. To better understand these experimental results, we shall study the Cabibbo-favored three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\omega\pi^{+}$ and $D^{0}\to K^{-}\omega\pi^{+}$ in this work, where the $\omega\pi$ pair is attributed to the decays of intermediate states $\rho(770)^{+}$ and $\rho(1450)^{+}$. The schematic diagram for the quasi-two body decay $D_{s}^{+}\to\eta\rho^{+}\to\eta\omega\pi^{+}$ is shown in Fig. 1. In the rest frame, the state $D_{s}^{+}$ will decay into the intermediate resonance $\rho^{+}$ along with the the bachelor state $\eta$, and then the $\rho^{+}$ decays into $\omega$ and $\pi^{+}$ within the strong interaction. The similar mechanisms can be occurred in $D^{+}$ and $D^{0}$ decay processes, with $\eta$ replaced by $K^{0}_{S}$ and $K^{-}$, respectively. The subprocesses $\rho(770,1450)^{+}\to\omega\pi^{+}$ in these decays will be introduced into their decay amplitudes in the isobar formalism pr135.B551 ; pr166.1731 ; prd11.3165 via the vector form factor $F_{\omega\pi}$. This form factor has been measured in the related processes of $\tau$ decay and $e^{+}e^{-}$ annihilation.
@@
The contributions of $\rho(770,1450)\to\omega\pi$ for the three-body decays $B\to\bar{D}^{(*)}\omega\pi$ were studied in Ref. JHEP2401-047 very recently. The $\rho$ family resonance contributions of kaon pair have been explored in Refs. prd101.111901 ; prd103.056021 ; prd103.016002 ; cpc46-053104 ; prd107.116023 ; prd110.056001 ; prd109.116009 and in Refs. prd85.092016 ; prd93.052018 ; prd103.114028 ; prd104.116019 for the three-body $B$ and $D$ meson decays, respectively. In Ref. 2309.00280 , four resonances $\rho(770)$, $\rho(1450)$, $\rho(1700)$ and $\rho(2150)$ have been employed to parametrize the related transition form factor for $\rho\to\omega\pi$ for the process $e^{+}e^{-}\to\omega\pi^{0}\to\pi^{+}\pi^{-}\pi^{0}\pi^{0}$ in the energy range $1.05$-$2.00$ GeV by SND Collaboration. But we will leave the contributions of $\rho(1700,2150)\to\omega\pi$ to future studies for the relevant decays, in view of that the masses of $\rho(1700,2150)$ are very close to or even beyond the masses of the initial $D^{+,0}$ and $D^{+}_{s}$ mesons, the contributions of $\omega\pi$ from $\rho(1700,2150)$ are unimportant comparing with that from $\rho(770)$ and $\rho(1450)$ 2309.00280 , and in addition, the excited $\rho$ states around $2$ GeV are not well understood plb813-136059 ; prd105.074035 .
On account of the $c$-quark mass, the heavy quark expansion tools and factorization approaches, which have been successfully used in hadronic $B$ meson decays for decades, encounter significant challenges when applied to the two-body or three-body hadronic $D$ meson decays. In this context, model independent methods, such as the topological-diagram approach prd85.054014 ; prd85.034036 ; prd86.014014 ; prd93.114010 ; prd100.093002 ; prd104.073003 ; prd110.094052 and factorization-assisted topological-amplitude approach prd86.036012 ; prd89.054006 ; adv-7627308 have been adopted in various $D$ decay studies. In Ref. prd104.116019 , we construct a theoretical framework for quasi-two-body $D$ meson decays with the help of electromagnetic form factors, with which we studied the contributions of $\rho(770,1450)\to K\bar{K}$ for the three-body $D$ decays within the flavour SU(3) symmetry. In this work we adopt the method in prd104.116019 and investigate the relevant decays within the quasi-two-body framework 1605.03889 ; prd96.113003 ; prd104.116019 , while neglect the interaction between $\omega\pi$ system and the corresponding bachelor state in relevant decay processes.
@@ II Differential branching fractions
where the Fermi coupling constant $G_{F}=1.1663788(6)\times 10^{-5}$ GeV-2 PDG-2024 , the product of the Cabbibo-Kobayashi-Maskawa (CKM) Cabibbo ; Kobayashi matrix elements $\lambda_{q}=V^{*}_{cq}V_{uq}$ and $\lambda_{b}=V^{*}_{cb}V_{ub}$, the $C$’s are Wilson coefficients at scale $\mu$, and the $O_{1}$ and $O_{2}$ are current-current operators, $O_{3}$-$O_{4}$ are QCD penguin operators, and $O_{8g}$ is chromomagnetic dipole operator. Then the total decay amplitude for the process of a $D$ meson decays into a pseudoscalar ($P$) plus a vector ($V$) can be described with typical topological diagram amplitudes $T_{P,V},C_{P,V},E_{P,V}$ and $A_{P,V}$ as well as additional penguin amplitudes in the factorization-assisted topological-amplitude approach prd89.054006 and the topological-diagram approach prd85.034036 ; prd93.114010 ; prd100.093002 ; prd104.073003 . One is referred to the Refs. prd89.054006 ; prd85.034036 ; prd93.114010 ; prd100.093002 ; prd104.073003 for the detailed discussions on these the topological diagram amplitudes.
The decay amplitude for $D^{0}\to K^{-}\rho(770)^{+}$ is dominated by the color-favored tree amplitude $T_{P}^{2B}$ with the $D^{0}\to K^{-}$ transition, where the superscript $2B$ stands for a two-body decay process. We have the amplitude $T_{P}^{2B}$ as prd89.054006 ; prd93.114010 ; prd100.093002
@@
where the subscript $\rho$ stands for the resonance $\rho(770)$ here, the $F_{1}^{D\to\bar{K}}(m^{2}_{\rho})$ is the transition form factor for $D^{0}\to K^{-}$ process. Beyond the dominated amplitude $T_{P}^{2B}$, there is a $W$-exchange nonfactorizable contribution from the amplitude $E^{2B}_{V}$, which can be parametrized as the form $E^{2B}_{V}=\mathcal{M}^{2B}_{E}\epsilon_{\rho}\!\cdot\!p_{D}$, and we have prd89.054006
@@
in the numerical calculations of the decays $D^{+}\to K^{0}_{S}\rho(770)^{+}$ and $D_{s}^{+}\to\eta\rho(770)^{+}$, respectively. With the parameters $\chi^{C}_{V}=-0.53$, $\phi^{C}_{V}=-0.25$, $\chi^{A}_{q}=0.11$ and $\phi^{A}_{q}=-0.35$ found in prd89.054006 .
@@
By connecting subprocess $\rho(770)^{+}\to\pi^{+}\pi^{0}$ and the two-body mode $D^{0}\to K^{-}\rho(770)^{+}$ together, we get the quasi-two-body decay $D^{0}\to K^{-}\rho(770)^{+}\to K^{-}\pi^{+}\pi^{0}$. Its decay amplitude can be written as prd104.116019
@@
where the summation is over the isovector resonances $\rho_{i}=\{\rho(770),\rho(1450),\rho(1700),...\}$ in $\rho$ family, with $m_{\rho_{i}}$ their masses. The $A_{i}$ and $\phi_{i}$ are the weights and phases for these resonances, respectively, and we can always have $A=1$ and $\phi=0$ for $\rho(770)$. Technically, the contributions from the excitations of $\omega(782)$ should also be include in Eq. (21), but their weights were found to be negligibly small jpg36-085008 .
@@
for the quasi-two-body decay $D^{0}\to K^{-}[\rho(770)^{+}\to]\omega\pi^{+}$ after the integration of $\cos\theta$. Then we can define a auxiliary electromagnetic form factor

### v2

@@ Abstract
Recently, the BESIII Collaboration has observed the three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\pi^{+}\omega$ and $D^{0}\to K^{-}\pi^{+}\omega$. In this work, we investigate the contributions of the subprocesses $\rho^{+}\to\omega\pi^{+}$ in these Cabibbo-favored decays $D\to h\omega\pi$, with $\rho^{+}=\{\rho(770)^{+},\rho(1450)^{+},\rho(770)^{+}\&\rho(1450)^{+}\}$ and $h=\{\eta,K^{0}_{S},K^{-}\}$, by introducing these subprocesses into the decay amplitudes of relevant decay processes via the vector form factor $F_{\omega\pi}$ which has measured in the related $\tau$ and $e^{+}e^{-}$ processes; we provide the first theoretical predictions for the branching fractions of the quasi-two-body decays $D_{s}^{+}\to\eta[\rho^{+}\to]\omega\pi^{+}$, $D^{+}\to K^{0}_{S}[\rho^{+}\to]\omega\pi^{+}$ and $D^{0}\to K^{-}[\rho^{+}\to]\omega\pi^{+}$. Our findings reveal that the contributions from the subprocess $\rho(770)^{+}\to\omega\pi^{+}$ are significant in these observed three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\omega\pi^{+}$ and $D^{0}\to K^{-}\omega\pi^{+}$, notwithstanding the contributions originating from the Breit-Wigner tail effect of $\rho(770)^{+}$. The numerical results of this study suggest that the dominant resonance contributions for the three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$ and $D^{+}\to K^{0}_{S}\omega\pi^{+}$ are originated from the $P$-wave intermediate states $\rho(770)^{+}$, $\rho(1450)^{+}$ and their interference effects.
@@ I Introduction
The three-body decay $D_{s}^{+}\to\eta\omega\pi^{+}$ has been observed for the first time by BESIII Collaboration recently, with its total branching fraction $\mathcal{B}=(0.54\pm 0.12_{\rm stat}\pm 0.04_{\rm syst})\%$ prd107.052010 . This numerical result aligns an earlier measurement by the CLEO Collaboration, which reported a branching fraction of $(0.85\pm 0.54_{\rm stat}\pm 0.06_{\rm syst})\%$ in Ref. prd80.051102 for the same decay channel, but the BESIII measurement offers significantly improved precision. Given the quark structure of the initial and the final states in this $D_{s}$ decay process, the Cabibbo-favored transition $c\to s$ along with the $W^{+}\to u\bar{d}$ will be the dominated process at quark level. The prospective intermediate states for this process could be the resonances $a_{0}(980)^{+}$, $\rho(770)^{+}$, $b_{1}(1235)^{+}$, and $\omega(1420)$, etc., and their excited states which will decay into the $\pi^{+}\eta$, $\omega\pi^{+}$ and $\omega\eta$ pairs in the final state, respectively PDG-2024 . The combination of $\pi^{+}$ and $\omega\eta$ with the intermediate resonance $\omega(1420)$ means a pure annihilation process for this $D_{s}^{+}$ decay. The union of $\omega$ and $a_{0}(980)^{+}\to\pi^{+}\eta$ for this decay is very similar to the decay process $D_{s}^{+}\to\pi^{0}[a_{0}(980)^{+}\to]\pi^{+}\eta$, which has been measured by BESIII prl123.112001 and the branching fraction was found to be consistent with the triangle rescattering processes via the intermediate processes $D_{s}^{+}\to\eta^{(\prime)}\rho^{+}$, $D_{s}^{+}\to K^{*}\bar{K}$ and $D_{s}^{+}\to K\bar{K^{*}}$ epjc80-895 ; plb803-135279 ; epjc80-1041 ; prd103.116016 ; prd109.076027 ; epjc81-1093 . But the branching fraction for $D_{s}^{+}\to\omega a_{0}(980)^{+}$ was found to be less than $3.4\times 10^{-4}$ in Ref. epjc81-1093 , due to the cancellation of the rescattering effects and the suppressed short-distance $W$ annihilation contribution. Consequently, the dominated contributions for the decay $D_{s}^{+}\to\eta\omega\pi^{+}$ are expected to come from the combination of $\eta$ and the resonances which will decay into $\omega\pi^{+}$ pair in the final state.
@@
The light meson pair of $\omega\pi$ which originating from the weak current of the matrix element $\langle\omega\pi|V^{\mu}\!-\!A^{\mu}|0\rangle$ in the three-body hadronic $D$ and $B$ meson decays is related to the processes $\tau\to\omega\pi\nu_{\tau}$ and $e^{+}e^{-}\to\omega\pi$. The $G$-parity conservation requires that the $\omega\pi$ pair is determined mainly by the vector part $V^{\mu}$ of the weak current zpc57-495 , which is associated with the $\rho$ family resonances as the intermediate states via the $P$-wave transition amplitudes prd86.037302 . The axial-vector part $A^{\mu}$ of the weak current for $\omega\pi$ system goes to the resonance $b_{1}(1235)$ and its excited states. The state $b_{1}(1235)$, with the quantum numbers $J^{PG}=1^{++}$, is related to the second-class weak current classified in terms of parity and $G$ parity pr112.1375 . The second-class weak current has been searched in $\tau$ decays in recent years, but no evidence has been observed ar0807.4900 ; npb218-110 . The decay amplitude for the $b_{1}(1235)\to\omega\pi$ subprocess through $S$- and $D$-waves in relevant processes is proportional to the mass difference between $u$ and $d$ quarks PDG-2024 , which makes the contribution of $\omega\pi$ from the axial-vector resonance $b_{1}(1235)$ accompanied by the second-class current highly suppressed.
@@
The ordinary decay mode for $\rho(770)\to\omega\pi$ is not allowed because of the pole mass of $\rho(770)$, which is apparently below the threshold of the $\omega\pi$ pair PDG-2024 . But the Breit-Wigner (BW) BW-model tail effect of $\rho(770)$, also known as the virtual contribution plb25-294 ; Dalitz62 ; prd94.072001 ; plb791-342 , was found indispensable to the production of $\omega\pi$ pair in the processes of $\tau\to\omega\pi\nu_{\tau}$ plb185-223 ; prd61.072003 ; rmp78.1043 ; prl103.041802 and $e^{+}e^{-}\to\omega\pi^{0}$ plb174-453 ; plb486-29 ; plb562-173 ; plb669-223 ; JETPL94-734 ; prd88.054013 ; prd94.112001 ; prd96.092009 ; plb813-136059 ; 2309.00280 . The excited state $\rho(1450)$ has been observed to decay into $\omega\pi$ pair in $B$ meson decays by the CLEO, Babar and Belle collaborations prd64.092001 ; prd74.012001 ; prd92.012013 . This state has been suggested as a $2S$-hybrid mixture in Ref. prd60.114011 in view of its decay characters npb443-233 ; prd52.1706 ; prd56.1584 , but its mass is consistent with the $2S$ excitation of $\rho(770)$ prd55.4157 . The further investigation of the interference between the $\rho(1450)$ and its ground state will provide deeper insights into its nature.
@@
In addition to data for the $D_{s}^{+}\to\eta\omega\pi^{+}$ decay, the absolute branching fractions of the decays $D^{0}\to K^{-}\pi^{+}\omega$, $D^{0}\to K^{0}_{S}\pi^{0}\omega$, and $D^{+}\to K^{0}_{S}\pi^{+}\omega$ have been determined by BESIII Collaboration recently in Ref. prd105.032009 , with the results to be $(3.392\pm 0.044_{\rm stat}\pm 0.085_{\rm syst})\%$, $(0.848\pm 0.046_{\rm stat}\pm 0.031_{\rm syst})\%$, and $(0.707\pm 0.041_{\rm stat}\pm 0.029_{\rm syst})\%$, respectively. To better understand these experimental results, we shall study the Cabibbo-favored three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\omega\pi^{+}$ and $D^{0}\to K^{-}\omega\pi^{+}$ in this work, where the $\omega\pi$ pair is attributed to the decays of intermediate states $\rho(770)^{+}$ and $\rho(1450)^{+}$. The schematic diagram for the quasi-two body decay $D_{s}^{+}\to\eta\rho^{+}\to\eta\omega\pi^{+}$ is shown in Fig. 1. In its rest frame, the state $D_{s}^{+}$ will decay into the intermediate resonance $\rho^{+}$ along with the the bachelor state $\eta$, and then the $\rho^{+}$ decays into $\omega$ and $\pi^{+}$ via the strong interaction. The similar pattern will arise in $D^{+}$ and $D^{0}$ decay processes, with the $\eta$ replaced by $K^{0}_{S}$ and $K^{-}$, respectively. The subprocesses $\rho(770,1450)^{+}\to\omega\pi^{+}$ in these decays will be introduced into their decay amplitudes in the isobar formalism pr135.B551 ; pr166.1731 ; prd11.3165 via the vector form factor $F_{\omega\pi}$. This form factor has been measured in the related processes of $\tau$ decay and $e^{+}e^{-}$ annihilation.
@@
The contributions of $\rho(770,1450)\to\omega\pi$ for the three-body decays $B\to\bar{D}^{(*)}\omega\pi$ were studied in Ref. JHEP2401-047 very recently. The $\rho$ family resonance contributions of kaon pair have been explored in Refs. prd101.111901 ; prd103.056021 ; prd103.016002 ; cpc46-053104 ; prd107.116023 ; prd110.056001 ; prd109.116009 and in Refs. prd85.092016 ; prd93.052018 ; prd103.114028 ; prd104.116019 for the three-body $B$ and $D$ meson decays, respectively. In Ref. 2309.00280 , four resonances $\rho(770)$, $\rho(1450)$, $\rho(1700)$ and $\rho(2150)$ have been employed to parametrize the related transition form factor for $\rho\to\omega\pi$ for the process $e^{+}e^{-}\to\omega\pi^{0}\to\pi^{+}\pi^{-}\pi^{0}\pi^{0}$ in the energy range 1.051.05-$2.00$ GeV by SND Collaboration. But we will leave the contributions of $\rho(1700,2150)\to\omega\pi$ to future studies for the relevant decays, in view of that the masses of $\rho(1700,2150)$ are very close to or even beyond the masses of the initial $D^{+,0}$ and $D^{+}_{s}$ mesons, the contributions of $\omega\pi$ from $\rho(1700,2150)$ are unimportant comparing with that from $\rho(770)$ and $\rho(1450)$ 2309.00280 , and in addition, the excited $\rho$ states around $2$ GeV are not well understood plb813-136059 ; prd105.074035 .
Due to the $c$-quark mass, the heavy quark expansion tools and the factorization approaches, which have been successfully used in hadronic $B$ meson decays for decades, encounter significant challenges when applied to the two-body or three-body hadronic $D$ meson decays. In this context, model independent methods, such as the topological-diagram approach prd85.054014 ; prd85.034036 ; prd86.014014 ; prd93.114010 ; prd100.093002 ; prd104.073003 ; prd110.094052 and factorization-assisted topological-amplitude approach prd86.036012 ; prd89.054006 ; adv-7627308 ; ar2503.18593 have been adopted in various $D$ decay studies. In Ref. prd104.116019 , we construct a theoretical framework for quasi-two-body $D$ meson decays with the help of electromagnetic form factors, with which we studied the contributions of $\rho(770,1450)\to K\bar{K}$ for the three-body $D$ decays within the flavour SU(3) symmetry. In this work we adopt the method in prd104.116019 and investigate the concerned decays within the quasi-two-body framework 1605.03889 ; prd96.113003 ; prd104.116019 , while neglect the interaction between $\omega\pi$ system and the corresponding bachelor state in relevant decay processes.
@@ II Differential branching fractions
where the Fermi coupling constant $G_{F}=1.1663788(6)\times 10^{-5}$ GeV-2 PDG-2024 , the product of the Cabbibo-Kobayashi-Maskawa (CKM) Cabibbo ; Kobayashi matrix elements $\lambda_{q}=V^{*}_{cq}V_{uq}$ and $\lambda_{b}=V^{*}_{cb}V_{ub}$, the $C$’s are Wilson coefficients at scale $\mu$, and the $O_{1}$ and $O_{2}$ are current-current operators, $O_{3}$-$O_{4}$ are QCD penguin operators, and $O_{8g}$ is chromomagnetic dipole operator. Then the total decay amplitude for the process of a $D$ meson decays into a pseudoscalar ($P$) plus a vector ($V$) can be described with typical topological diagram amplitudes $T_{P,V},C_{P,V},E_{P,V}$ and $A_{P,V}$ according to the diagrams shown in Fig. 2, as well as additional penguin amplitudes in the factorization-assisted topological-amplitude approach prd89.054006 and the topological-diagram approach prd85.034036 ; prd93.114010 ; prd100.093002 ; prd104.073003 . One is referred to the Refs. prd89.054006 ; prd85.034036 ; prd93.114010 ; prd100.093002 ; prd104.073003 for the detailed discussions on these the topological diagram amplitudes.
Figure 2: Typical topological diagrams for the concerned decays at quark level, where $\rho$ stands for the intermediate states $\rho(770,1450)^{+}$, $h$ for $\eta,K^{0}_{S}$ or $K^{-}$, and the symbol $\otimes$ stands for the weak vertex in this work.
The decay amplitude for $D^{0}\to K^{-}\rho(770)^{+}$ is dominated by the color-favored tree amplitude $T_{P}^{2B}$ with the $D^{0}\to K^{-}$ transition as shown in Fig. 2-(a), where the superscript $2B$ stands for a two-body decay process. We have the amplitude $T_{P}^{2B}$ as prd89.054006 ; prd93.114010 ; prd100.093002
@@
where the subscript $\rho$ stands for the resonance $\rho(770)$ here, the $F_{1}^{D\to\bar{K}}(m^{2}_{\rho})$ is the transition form factor for $D^{0}\to K^{-}$ process. Beyond the dominated amplitude $T_{P}^{2B}$, there is a $W$-exchange nonfactorizable contribution from the amplitude $E^{2B}_{V}$ from the topological diagram Fig. 2-(c), which can be parametrized as the form $E^{2B}_{V}=\mathcal{M}^{2B}_{E}\epsilon_{\rho}\!\cdot\!p_{D}$, and we have prd89.054006
@@
in the numerical calculations of the decays $D^{+}\to K^{0}_{S}\rho(770)^{+}$ and $D_{s}^{+}\to\eta\rho(770)^{+}$ according to the topological diagrams Fig. 2-(b) and -(d), respectively. With the parameters $\chi^{C}_{V}=-0.53$, $\phi^{C}_{V}=-0.25$, $\chi^{A}_{q}=0.11$ and $\phi^{A}_{q}=-0.35$ found in prd89.054006 .
@@
By connecting subprocess $\rho(770)^{+}\to\pi^{+}\pi^{0}$ and the two-body mode $D^{0}\to K^{-}\rho(770)^{+}$ together, we will get the quasi-two-body decay $D^{0}\to K^{-}\rho(770)^{+}\to K^{-}\pi^{+}\pi^{0}$. Its decay amplitude can be written as prd104.116019
@@
where the summation is over the isovector resonances $\rho_{i}=\{\rho(770),\rho(1450),\rho(1700),$...\} in $\rho$ family, with $m_{\rho_{i}}$ their masses. The $A_{i}$ and $\phi_{i}$ are the weights and phases for these resonances, respectively, and we can assign $A=1$ and $\phi=0$ for $\rho(770)$. Technically, the contributions from the excitations of $\omega(782)$ should also be include in Eq. (21), but their weights were found to be negligibly small jpg36-085008 .
@@
for the quasi-two-body decay $D^{0}\to K^{-}[\rho(770)^{+}\to]\omega\pi^{+}$ after the integration of $\cos\theta$. Then we can define an auxiliary electromagnetic form factor